A train is approaching a signal tower at a speed of 40 m/s. The train engineer sounds the 1000 Hz whistle and a switchman inA train is approaching a signal tower at a speed of 40 m/s. The train engineer sounds the 1000 Hz whistle and a switchman in the tower responds by sounding the 1200 Hz siren. The air is still and the speed of sound is 340 m/s. 9) In Situation 16.4, the wavelength of the train whistle tone reaching the switchman, in SI units, is given by:

A) 0.32

B) 0.34

C) 0.30

D) 0.36

E) 0.38
Answer #1

considering doppler effect

f = fo v / ( v - vs)

f = 1000* 340 / ( 340 - 40)

f = 1130 Hz

L = c/f = 340 / 1130

L= 0.30